# Robust Signal Detection with Quadratically Convex Orthosymmetric Constraints

### v2

@@ 5 Discussion
Similarly, the upper bounds $E_{\text{alt}}^{2}(\epsilon,K,N,\sigma^{2})$ and $P_{\text{alt}}^{2}(\epsilon,K,N,\sigma^{2})$ are obtained by substituting all $k_{1}^{\star}$ and $k_{2}^{\star}$ in $E^{2}(\epsilon,K,N,\sigma^{2})$ and $P^{2}(\epsilon,K,N,\sigma^{2})$ by $k_{\text{alt}}^{\star}$. With the proof in Appendix B.7, one can verify that $k_{\text{alt}}^{\star},P_{\text{alt}}^{\star}$ and its corresponding testing process are equivalent to that of $k_{1}^{\star}$, $k_{2}^{\star}$, $P_{1}^{\star}$, and $P_{2}^{\star}$. The new lower bound (5.1) and the upper bounds $E_{\text{alt}}^{2}(\epsilon,K,N,\sigma^{2})$ and $P^{2}(\epsilon,K,N,\sigma^{2})$ do match the previous lower bound (3.6) and the upper bounds (3.3), (3.5) as well.
The results developed in Section 3 should be compared and contrasted with the existing results for the corresponding estimation problem. Specifically, in the estimation setting — where the same distribution assumptions and corruption model are imposed, but the goal is to estimate the mean vector $\mu$ — the minimax rates are established by Donoho et al., (1990) and Neykov, (2022). Notably, the Kolmogorov widths and the associated optimal dimensions again play a central role in these results.
@@
We refer readers to Appendix B.11 for a straightforward computational verification of Lemma 5.3. Following the discussion after Corollary 3.12, we can also analyze the phase transition of the estimation problem. It occurs around $\epsilon\asymp\sqrt{\frac{k_{e}^{\star}}{N}}$. If $\epsilon\lesssim\sqrt{\frac{k_{e}^{\star}}{N}}$, then $\epsilon$ is too small to influence the minimax rate of the estimation error. Otherwise (for example, $\epsilon\asymp N^{-\frac{1}{4}}$), the adversary is strong enough to the extent that $N$ is excluded and the minimax rate only relies on $\epsilon$ and $\sigma$.
@@ 5.1 Open Problems and Future Work
In other words, rather than being fixed at the origin, the mean vector under the null is allowed to vary within a neighborhood of the origin. An interesting corollary of their work is that as the neighborhood expands, the minimax testing rate converges to the minimax estimation rate — an effect that parallels a key phenomenon in our setting. Given the framework of their problem, it is natural to ask whether one can establish the minimax rate with constrained and adversarial setting of (5.5). From the perspective of the testing-estimation relationship, an even more important question is: under what conditions does the gap between testing and estimation disappear?
@@ Acknowledgments
Appendix A Some Basic Probabilistic Concepts and Useful Lemmas
@@
Definition A.1 (Total variation distance).
@@
$\text{TV}(\mathcal{P},\mathcal{Q})=\displaystyle\int_{\Omega_{\nu}}\min\limits\left\{p,q\right\}d\nu=\frac{1}{2}\displaystyle\int_{\Omega_{\nu}}\left|p-q\right|d\nu.$ (A.1)
Definition A.2 (Kullback-Leibler divergence (KL divergence)).
@@
Lemma A.3 (Pinsker’s inequality).
@@
$\text{TV}(\mathcal{P},\mathcal{Q})\leq\sqrt{2\text{KL}(\mathcal{P},\mathcal{Q})}.$
@@
$\text{TV}(\mathcal{P},\mathcal{Q})\leq\sqrt{2\min\limits\left\{\text{KL}(\mathcal{P},\mathcal{Q}),\text{KL}(\mathcal{Q},\mathcal{P})\right\}}.$
@@
Definition A.6 (sub-Gaussian random variables).
@@
Definition A.7 (sub-exponential random variables).
@@
Lemma A.8 (Bernstein’s inequality).
@@
$\mathbb{P}\left(\sum\limits_{i=1}^{n}X_{i}\geq t\right)\leq\exp\left\{-ct^{2}/\sum\limits_{i=1}^{n}\left\lVert X_{i}\right\rVert_{\psi_{2}}^{2}\right\}.$ (A.3)
@@
$\mathbb{P}\left(\sum\limits_{i=1}^{n}X_{i}\geq t\right)\leq\exp\left\{-c\min\limits\left\{t^{2}/\sum\limits_{i=1}^{n}\left\lVert X_{i}\right\rVert_{\psi_{1}}^{2},t/\max\limits_{i}\left\{\left\lVert X_{i}\right\rVert_{\psi_{1}}\right\}\right\}\right\}.$ (A.4)
@@
The famous Hanson-Wright inequality is a powerful tool to bound the tail probability of quadratic forms of the sub-Gaussian random variables. It can also be viewed as a generalization of the Bernstein’s inequality (A.4). See Hanson and Wright, (1971) for the initial paper and Rudelson and Vershynin, (2013) for a modern proof.
Lemma A.10 (The Hanson-Wright inequality).
@@
$\mathbb{P}\left(|X^{\top}AX-\mathbb{E}X^{\top}AX|>t\right)\leq 2\exp\left\{-c\min\limits\left\{\frac{t^{2}}{K^{4}\left\lVert A\right\rVert_{F}^{2}},\frac{t}{K^{2}\left\lVert A\right\rVert_{2}}\right\}\right\},$ (A.5)
@@
Definition A.11 (The Wishart matrix).
@@
We can leverage Lemma A.12 to bound the $\ell_{2}$ operator norm of $W_{d}(n)$.
Corollary A.13.
@@
Set $\bm{\Sigma}=\mathbf{I}_{d}$ in the first inequality of Lemma A.12, we have:
@@
By Corollary A.13, we already have $\left\lVert\sum\limits_{i=1}^{n}Z_{i}Z_{i}^{\top}-n\mathbf{I}_{d}\right\rVert_{2}\leq C_{1}\left(\sqrt{nd}+\sqrt{n\ln\left(\frac{1}{\delta}\right)}+\ln\left(\frac{1}{\delta}\right)\right)$. We also have $n\left\lVert\mu\mu^{\top}\right\rVert_{2}=n\alpha^{2}$. It remains to bound $\left\lVert\sum\limits_{i=1}^{n}\mu Z_{i}^{\top}\right\rVert_{2}$.
@@
Since $Z_{i}\sim\mathcal{N}(0,\mathbf{I}_{d})$ for each $i$, from A.9 we know that $\left\lVert\sum\limits_{i=1}^{n}\frac{Z_{i}}{\sqrt{n}}-d\right\rVert_{2}$ is a sub-Gaussian random variable and therefore with probability higher than $1-\delta$ we have:
@@
The next lemma is from Fact 4.2 in Dong et al., (2019). It is an analog of Corollary A.13, but within the field when the $\ell_{1}$ norm of a weight $\omega$ is small compared to $n$.
Lemma A.15.
@@
Corollary A.16.
Assuming the same conditions as Lemma A.15, we further allow $X_{i}$ from $\mathcal{N}(\mu,\mathbf{I}_{d})$ with $\left\lVert\mu\right\rVert_{2}\leq\alpha$. As a consequence, the inequality will track the dependence on $\alpha$:
@@
The proof is basically from Lemma A.15 the standard argument of concentration. For any $v$ such that $\left\lVert v\right\rVert_{2}=1$, we have:
@@
Bound of the first term can be obtained from Lemma A.15. For the second term, by the Cauchy-Schwarz inequality we can bound it by $2\left\lVert\omega\right\rVert_{1}\cdot\alpha^{2}\leq C\cdot\epsilon n\alpha^{2}$. For the third term, since $\sum\limits_{i=1}^{n}\omega_{i}v^{\top}(X_{i}-\mu)$ follows from $\mathcal{N}(0,\left\lVert\omega\right\rVert_{2}^{2})$, we know we can bound it by $C\cdot\alpha\left\lVert\omega\right\rVert_{2}\leq C\cdot\alpha\sqrt{\epsilon n}$. Combining them together yields the result. ∎
@@
Lemma A.17 (Hölder’s inequality).
@@
Finally, we introduce an intuitive and useful result in convexity and combinatorial mathematics. We note that Canonne et al., (2023) first introduced this result in their paper but their proof contains a flaw which we fix below.
Lemma A.19.
@@
Appendix B Deferred Proofs of the Lemmas and Main Theorems
B.1 Proof of Lemma 3.1
The proof of Lemma 3.1 leverages key properties of QCO sets and applies the minimax theorem to a suitably truncated vector. A similar argument appears in Section 3.3.2 of Neykov, (2022).

### v3

@@ 6.2 Equivalence in Optimal Dimensions
Similarly, the upper bounds $E_{\text{alt}}^{2}(\epsilon,K,N,\sigma^{2})$ and $P_{\text{alt}}^{2}(\epsilon,K,N,\sigma^{2})$ are obtained by substituting all $k_{1}^{\star}$ and $k_{2}^{\star}$ in $E^{2}(\epsilon,K,N,\sigma^{2})$ and $P^{2}(\epsilon,K,N,\sigma^{2})$ by $k_{\text{alt}}^{\star}$. With the proof in Appendix C.7, one can verify that $k_{\text{alt}}^{\star},P_{\text{alt}}^{\star}$ and its corresponding testing process are equivalent to that of $k_{1}^{\star}$, $k_{2}^{\star}$, $P_{1}^{\star}$, and $P_{2}^{\star}$. The new lower bound (6.1) and the upper bounds $E_{\text{alt}}^{2}(\epsilon,K,N,\sigma^{2})$ and $P^{2}(\epsilon,K,N,\sigma^{2})$ do match the previous lower bound (3.8) and the upper bounds (3.5), (3.7) as well.
6.3 Testing-Estimation Relationship
The results developed in Section 3 should be compared and contrasted with the existing results for the corresponding estimation problem. Specifically, in the estimation setting — where the same distribution assumptions and corruption model are imposed, but the goal is to estimate the mean vector $\mu$ — the minimax rates are established by Donoho et al., (1990) and Neykov, (2022). Although the authors do not explicitly employ the concepts of Kolmogorov widths and optimal dimensions, their results can nonetheless be interpreted within our framework. Notably, the Kolmogorov widths and the associated optimal dimensions again play a central role in these results.
@@ 6.3 Testing-Estimation Relationship
We refer readers to Appendix C.11 for a straightforward computational verification of Lemma 6.3. Following the discussion after Corollary 3.15, we can also analyze the phase transition of the estimation problem. It occurs around $\epsilon\asymp\sqrt{\frac{k_{e}^{\star}}{N}}$. If $\epsilon\lesssim\sqrt{\frac{k_{e}^{\star}}{N}}$, then $\epsilon$ is too small to influence the minimax rate of the estimation error. Otherwise (for example, $\epsilon\asymp N^{-\frac{1}{4}}$), the adversary is strong enough to the extent that $N$ is excluded and the minimax rate only relies on $\epsilon$ and $\sigma$.
@@ 6.4 Open Problems and Future Work
In other words, rather than being fixed at the origin, the mean vector under the null is allowed to vary within a neighborhood of the origin. An interesting corollary of their work is that as the neighborhood expands, the minimax testing rate converges to the minimax estimation rate — an effect that parallels a key phenomenon in our setting. Such match in estimation and testing rate is also observed by other work (e.g., Ndaoud, (2019)). Given the framework of their problem, it is natural to ask whether one can establish the minimax rate with constrained and adversarial setting of (6.5). From the perspective of the testing-estimation relationship, an even more important question is: under what conditions does the gap between testing and estimation disappear?
@@ Acknowledgments
Appendix A Theoretical and Polynomial-Time Algorithms
A.1 Theoretical Algorithm
According to Theorem 3.8 and Section C.5, we first attempt to search for a consistent set $T$ with respect to $\phi_{e}$. $\phi_{e}(\mathbf{X}_{T})$ then determines whether we are to accept or reject $H_{0}$.
1
2import $\mathbf{X},N,K,\sigma$
3
4Compute $k_{1}^{\star},P_{1}^{\star},k_{2}^{\star},P_{2}^{\star}$ from $N,K,\sigma,\epsilon$ as Definition 2.3 and 2.4
5if $k_{1}^{\star}\leq k_{2}^{\star}$ then
6 $k^{\star}=k_{1}^{\star},P^{\star}=P_{1}^{\star}$
7
8else
9 $k^{\star}=k_{2}^{\star},P^{\star}=P_{2}^{\star}$
10
[…]
13
14 consistent_flag == True
16    subset_test == True
17 else
18    subset_test == False
[…]
20
22       subsubset_test == True
23    else
24       subsubset_test == False
25    if subset_test XOR subsubset_test then
26       consistent_flag == False
27       break
28
29 if consistent_flag then
30    if subset_test then
31         Reject $H_{0}$
32       return
33    else
34         Accept $H_{0}$
35       return
36
37
38Accept $H_{0}$
return
Algorithm 3 theoretical algorithm when $\rho\gtrsim E(\epsilon,K,N,\sigma^{2})$
A.2 Polynomial-Time Algorithm
The polynomial-time algorithm combines the three sub-algorithms in Section C.6 together to filter the observations $\mathbf{X}$ and through the weight $\omega$. The theoretical guarantee is provided in Section C.6.
2while $\texttt{i}<N$ do
4
5    $\texttt{count}=\texttt{count}+1$
6    if $\texttt{count}>\epsilon N$ then
7       return None
8     Delete $Y_{i}$ from $\mathbf{Y}$
9 $\texttt{i}=\texttt{i}+1$
10
return $\mathbf{Y}$
Algorithm 4 Prefiltering.
1Set $\gamma_{2}$ as (C.13), and $\lambda=\left\lVert Y^{\top}D(\omega)Y-N\mathbf{I}_{k^{\star}}\right\rVert_{2}$. ($\omega$ is initialized as $\mathbf{1}$.)
2while $\lambda\geq\gamma_{2}$ do
3   Set $v$ to be the unit singular vector associated with $\lambda$
6  Update $w$ according to (C.14)
7 if $\left\lVert\omega\right\rVert_{1}<N(1-2\epsilon)$ then
8
9    return None
10
12
return $\omega$
Algorithm 5 Sample filtering when $N>k^{\star}$.
1Set $\gamma_{3}$ as (C.21), and $\lambda=\left\lVert\sqrt{D(\omega)}\mathbf{Y}\mathbf{Y}^{\top}\sqrt{D(\omega)}-D(\omega)\right\rVert_{2}$. ($\omega$ is initialized as $\mathbf{1}$.)
2while $\lambda\geq\gamma_{3}$ do
3   Set $v$ to be the unit singular vector associated with $\lambda$
5  Update $w$ according to (C.22)
6 if $\left\lVert\omega\right\rVert_{1}<N(1-6\epsilon)$ then
7
8    return None
9
11
return $\omega$
Algorithm 6 Sample filtering when $N\leq k^{\star}$.
3Set $\omega_{i_{1}},\omega_{i_{2}},\dots,\omega_{i_{\epsilon N}}$ to zero
return $\omega$
Algorithm 7 Weight filtering.
1
2import $\mathbf{X},N,K,\sigma=1$
3import Algorithm 4, 5, 6, 7
4
[…]
6Compute $k_{1}^{\star},P_{1}^{\star},k_{2}^{\star},P_{2}^{\star}$ from $N,K,\sigma=1,\epsilon$ as Definition 2.3 and 2.4
7if $k_{1}^{\star}\leq k_{2}^{\star}$ then
8 $k^{\star}=k_{1}^{\star},P^{\star}=P_{1}^{\star}$
9
10else
11 $k^{\star}=k_{2}^{\star},P^{\star}=P_{2}^{\star}$
12
[…]
14Execute Algorithm 4, and record the returning value $\texttt{R}_{1}$
15if $\texttt{R}_{1}$ is None then
16   Reject $H_{0}$
17 return
18
19else
20 $\mathbf{Y}=\texttt{R}_{1}$
21
22Set $\omega=\mathbf{1}_{[k^{\star}]}$
23if $N>k^{\star}$ then
24   Execute Algorithm 5, and record the returning value $\texttt{R}_{2}$
25else
26   Execute Algorithm 6, and record the returning value $\texttt{R}_{2}$
27if $\texttt{R}_{2}$ is None then
28   Reject $H_{0}$
29 return
30
31else
32 $\omega=\texttt{R}_{2}$
33
34Execute Algorithm 7, and record the returning value $\texttt{R}_{3}$
[…]
37   Reject $H_{0}$
38
39else
40   Accept $H_{0}$
41
return
Algorithm 8 polynomial-time algorithm when $\rho\gtrsim P(\epsilon,k^{\star},N,1)$.
A.3 Deferred Experiment Results
The following figures display additional results deferred from Section 5.4.
Figure 4: Top: Rejection rate (or Empirical power) versus the $\ell_{2}$ norm of the mean under $H_{1}$ in the classic setting ($N=2000$). Bottom: Rejection rate versus the $\ell_{2}$ norm of the mean in the high-dimensional setting ($N=500$).
Appendix B Some Basic Probabilistic Concepts and Useful Lemmas
@@
Definition B.1 (Total variation distance).
@@
$\text{TV}(\mathcal{P},\mathcal{Q})=\displaystyle\int_{\Omega_{\nu}}\min\limits\left\{p,q\right\}d\nu=\frac{1}{2}\displaystyle\int_{\Omega_{\nu}}\left|p-q\right|d\nu.$ (B.1)
Definition B.2 (the Kullback-Leibler divergence (KL divergence)).
@@
Lemma B.3 (Pinsker’s inequality).
@@
$\text{TV}(\mathcal{P},\mathcal{Q})\leq\sqrt{\frac{1}{2}\text{KL}(\mathcal{P},\mathcal{Q})}.$
@@
$\text{TV}(\mathcal{P},\mathcal{Q})\leq\sqrt{\frac{1}{2}\min\limits\left\{\text{KL}(\mathcal{P},\mathcal{Q}),\text{KL}(\mathcal{Q},\mathcal{P})\right\}}.$
@@
The coupling that achieves the equality above is referred as the optimal coupling between $\mathcal{P}$ and $\mathcal{Q}$.
@@
Definition B.6 (Sub-Gaussian random variables).
@@
Definition B.7 (Sub-exponential random variables).
@@
Lemma B.8 (Bernstein’s inequality).
@@
$\mathbb{P}\left(\sum\limits_{i=1}^{n}X_{i}\geq t\right)\leq\exp\left\{-ct^{2}/\sum\limits_{i=1}^{n}\left\lVert X_{i}\right\rVert_{\psi_{2}}^{2}\right\}.$ (B.3)
@@
$\mathbb{P}\left(\sum\limits_{i=1}^{n}X_{i}\geq t\right)\leq\exp\left\{-c\min\limits\left\{t^{2}/\sum\limits_{i=1}^{n}\left\lVert X_{i}\right\rVert_{\psi_{1}}^{2},t/\max\limits_{i}\left\{\left\lVert X_{i}\right\rVert_{\psi_{1}}\right\}\right\}\right\}.$ (B.4)
@@
The famous Hanson–Wright inequality is a powerful tool to bound the tail probability of quadratic forms of the sub-Gaussian random variables. It can also be viewed as a generalization of the Bernstein’s inequality (B.4). See Hanson and Wright, (1971) for the initial paper and Rudelson and Vershynin, (2013) for a modern proof.
Lemma B.10 (Hanson–Wright inequality).
@@
$\mathbb{P}\left(|X^{\top}AX-\mathbb{E}X^{\top}AX|>t\right)\leq 2\exp\left\{-c\min\limits\left\{\frac{t^{2}}{K^{4}\left\lVert A\right\rVert_{F}^{2}},\frac{t}{K^{2}\left\lVert A\right\rVert_{2}}\right\}\right\},$ (B.5)
@@
Definition B.11 (Wishart matrix).
@@
We can leverage Lemma B.12 to bound the $\ell_{2}$ operator norm of $W_{d}(n)$.
Corollary B.13.
@@
Set $\bm{\Sigma}=\mathbf{I}_{d}$ in the first inequality of Lemma B.12, we have:
@@
By Corollary B.13, we already have $\left\lVert\sum\limits_{i=1}^{n}Z_{i}Z_{i}^{\top}-n\mathbf{I}_{d}\right\rVert_{2}\leq C_{1}\left(\sqrt{nd}+\sqrt{n\ln\left(\frac{1}{\delta}\right)}+\ln\left(\frac{1}{\delta}\right)\right)$. We also have $n\left\lVert\mu\mu^{\top}\right\rVert_{2}=n\alpha^{2}$. It remains to bound $\left\lVert\sum\limits_{i=1}^{n}\mu Z_{i}^{\top}\right\rVert_{2}$.
@@
Since $Z_{i}\sim\mathcal{N}(0,\mathbf{I}_{d})$ for each $i$, from B.9 we know that $\left\lVert\sum\limits_{i=1}^{n}\frac{Z_{i}}{\sqrt{n}}-d\right\rVert_{2}$ is a sub-Gaussian random variable and therefore with probability higher than $1-\delta$ we have:
@@
The next lemma is from Fact 4.2 in Dong et al., (2019). It is an analog of Corollary B.13, but within the field when the $\ell_{1}$ norm of a weight $\omega$ is small compared to $n$.
Lemma B.15.
@@
Corollary B.16.
Assuming the same conditions as Lemma B.15, we further allow $X_{i}$ from $\mathcal{N}(\mu,\mathbf{I}_{d})$ with $\left\lVert\mu\right\rVert_{2}\leq\alpha$. As a consequence, the inequality will track the dependence on $\alpha$:
@@
The proof is basically from Lemma B.15 the standard argument of concentration. For any $v$ such that $\left\lVert v\right\rVert_{2}=1$, we have:
@@
Bound of the first term can be obtained from Lemma B.15. For the second term, by the Cauchy–Schwarz inequality we can bound it by $2\left\lVert\omega\right\rVert_{1}\cdot\alpha^{2}\leq C\cdot\epsilon n\alpha^{2}$. For the third term, since $\sum\limits_{i=1}^{n}\omega_{i}v^{\top}(X_{i}-\mu)$ follows from $\mathcal{N}(0,\left\lVert\omega\right\rVert_{2}^{2})$, we know we can bound it by $C\cdot\alpha\left\lVert\omega\right\rVert_{2}\leq C\cdot\alpha\sqrt{\epsilon n}$. Combining them together yields the result. ∎
@@
Lemma B.17 (Hölder’s inequality).
@@
Finally, we present an intuitive and useful result at the intersection of convexity and combinatorics. While this result was originally introduced in Canonne et al., (2023), their proof contains a technical gap which we address below.
Lemma B.19.
@@
Appendix C Deferred Proofs of the Lemmas and Main Theorems
C.1 Proof of Lemma 3.1